An electron, with a mass of 9.1 x 10–31 kg and charge of 1.6 x 10–19 C, is
accelerated to a velocity of 4.0 x 106
m/s then enters a uniform magnetic field
of 5.0 x 10–3 T at an angle of 90° to the direction of travel.
a. What is the radius of the circular path it follows?
b. Through what potential difference was the electron accelerated?
1
Expert's answer
2020-06-04T09:57:16-0400

As the electron moves in the magnetic field, two forces of equal magnitude are acting on it:


"qvB=\frac{mv^2}{R},\\
\space\\
R=\frac{mv}{qB}=4.55\text{ mm}."

Apply the work-energy theorem to find the potential difference:


"\frac{mv^2}{2}=q\phi,\\
\space\\
\phi=\frac{mv^2}{2q}=45.5\text{ V}."